can someone plss help me im very stuck ?
decide whether the given ordered pair is a solution of the given system.
1. (3,-7)
5x + y = 8
2x - 3y = 28


Answer by Edwin McCravy(20059)

    (3,-7)

ì5x +  y =  8
í
î2x - 3y = 28

    
Substitute the first number, 3, for x and the second number -7 for y
in both equations and see if you end up with the same number on both
sides of the equal sign.

Substituting 3 for x and -7 for y in the first equation of the system:

    5x +  y = 8
5(3) + (-7) = 8
     15 - 7 = 8
          8 = 8

So (3,-7) is a solution to the first equation.  But in order to be
a solution for the whole system, (3,-7) must also be a solution
to the second equation:

Substituting 3 for x and -7 for y in the second equation:

     2x - 3y = 28
2(3) - 3(-7) = 28
      6 + 21 = 28
          27 = 28

So (3,-7) is not a solution to the second equation so it is not
a solution to the system.

Edwin